OpenStudy (anonymous):

f(x)=-3*(1.5)^x-2.5 horizontal asymptote

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

y=-2.5

OpenStudy (anonymous):

y=1.5

OpenStudy (anonymous):

y=2.5

OpenStudy (michele_laino):

we have to evaluate this limit: \[\mathop {\lim }\limits_{x \to - \infty } \left\{ { - 3{{\left( {\frac{3}{2}} \right)}^x} - \frac{5}{2}} \right\} = ...?\]

OpenStudy (anonymous):

so the answer is -2.5

OpenStudy (michele_laino):

yes!